# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

### v1

@@ 4 Outlook and conclusions
The new approach mainly consists in splitting the calculation of matrix elements, which had been so far performed using a single large GPU kernel, into several smaller kernels. The first optimization, which parallelizes the calculation for different helicities of the initial and final state particles to different CUDA Streams, achieves a reduction by one to two orders of magnitude in the number of events that must be computed in parallel to make an efficient use of the GPU. This is interesting from a user perspective, as it should allow event generation jobs on GPUs using smaller numbers of events than the current CUDACPP. The second and third optimizations reorganize the calculation of color sums as a separate GPU kernel, which is taken as the default implementation, and possibly as a cuBLAS calculation, which can be enabled at runtime. Delegating the color sum to a separate kernel approximately provides a 10% to 20% improvement in peak throughputs. The cuBLAS implementation has a worse performance in most, but not all, configurations, but it is very interesting as it allows the use of NVidia Tensor Cores, while the existing CUDACPP kernels only use traditional CUDA cores. The fourth development consists in splitting up the computation of different Feynman diagrams in different GPU kernels: its performance is poor, mainly because of kernel launch overheads, but this may represent a useful starting point for further developments, possibly using CUDA Graphs.
@@ Acknowledgements
I thank the MG5aMC and the Madgraph on GPU development teams for our fruitful collaboration on these projects. In particular, I warmly thank Olivier Mattelaer for our initial discussions and our coding together for the new developments shown in this paper during the 2022 GPU Hackathon and more generally for our effective collaboration over many years on the Madgraph on GPU developments. This project would not have been possible without him. I am grateful to the organisers and our mentors at the CSCS GPU Hackathon in Lugano in September 2022, and I thank the other participants from our team, Stephan Hageböck, Stefan Roiser and Zenny Wettersten, for our useful discussions. I kindly acknowledge the use of LUMI HPC resources under project 465001592 (CERN / Madgraph GPU porting, EHPC-DEV-2024D11-007) to derive the results I showed for AMD GPUs.

### v2

@@ 5 Summary and outlook
The first large batch of changes, which includes three sets of optimizations, has already been merged upstream and integrated in a new production release of the CUDACPP software bib:cudacpp364 , in collaboration with my colleagues. The first optimization, which parallelizes the calculation for different helicities of the initial and final state particles to different CUDA Streams, achieves a reduction by one to two orders of magnitude in the number of events that must be computed in parallel to make an efficient use of the GPU. This is interesting from a user perspective, as it should allow event generation jobs on GPUs using smaller numbers of events than the current CUDACPP. The second and third optimizations reorganize the calculation of color sums as a separate GPU kernel, which is taken as the default implementation, and also offer the possibly to perform it using GPU-specific BLAS libraries. On NVidia GPUs, in particular, the cuBLAS implementation has been optimized to a level where it is faster than kernels for moderately complex processes such as $gg\!\rightarrow\!t\bar{t}ggg$; this is also interesting as it may offer a mechanism for CUDACPP to exploit tensor cores on NVidia GPUs, which are otherwise unused.
The second large batch of enhancements, which includes three further sets of software changes, was still work in progress at the time of my previous preprint bib:preprintV1 of this paper, but it has now been completed and I also recommend its inclusion in a new production release of the software. It is available as a github tag bib:ihel6tag , and a pull request for its merge upstream has been created. These three sets of changes all consist in different ways to split up the computation of Feynman diagrams into separate GPU kernels. The fourth development, which uses a different GPU kernel for each Feynman diagram, has a very poor performance, even when the workflow is orchestrated by CUDA graphs, because it implies very heavy access to GPU global memory. The fifth development allows the definition of kernels that compute larger groups of Feynman diagrams, possibly defined in separate source code files. The sixth development goes back to a single GPU kernel for all Feynman diagrams, which however delegates their calculation to different device functions for separate diagram groups. All these changes are fully functional and have been extensively tested on all platforms.
While this second line of R&D started off as an attempt to further improve throughputs for our standard candles such as $gg\!\rightarrow\!t\bar{t}ggg$, it now seems clear to me that these processes are best computed using the traditional CUDACPP approach of a single GPU kernel for all Feynman diagrams, since they are simple enough to afford this: the same approach and the same throughputs are still achievable using the new software, which can be configured to produce a single diagram group. The interest of the new feature, conversely, is in much more complex physics calculations such as those required by $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes. I have shown, in particular, that diagram splitting makes it possible to compute matrix elements for $gg\!\rightarrow\!t\bar{t}gggg$ in both the GPU and vector CPU implementations of CUDACPP, and for $gg\!\rightarrow\!t\bar{t}ggggg$ in the latter. The code I used for the $gg\!\rightarrow\!t\bar{t}gggg$ tests that I presented is also available as a github tag bib:ihel6tag4g , which I created only for documentation purposes and not as a patch to integrate upstream.
Further optimizations of the internal CUDACPP engine for computing matrix elements, possibly extending the work I presented here, are of course possible. The clean separation of Feynman diagrams and color sums in this new implementation, for instance, makes it possible to imagine scenarios where the former is performed on vector CPUs and the latter on GPUs using cuBLAS for particularly complex processes. In my opinion, however, the ME calculation for LO processes in CUDACPP has now been optimized to a level where there are no obvious opportunities for further significant throughput speedups. Other areas of work have higher priority, such as in particular the integration of LO CUDACPP into the upcoming Madgraph7 with an enhanced phase space sampling based on MadNIS bib:madnis , or the extension of the CUDACPP infrastructure to NLO physics processes bib:zwnlo . The work I presented in this paper will be relevant in both of these activities.
@@ Acknowledgements
I thank the MG5aMC and the Madgraph on GPU development teams for our fruitful collaboration on these projects. In particular, I warmly thank Olivier Mattelaer for our initial discussions and our coding together for the new developments shown in this paper during the 2022 GPU Hackathon and more generally for our effective collaboration over many years on the Madgraph on GPU developments. This project would not have been possible without him. I thank Daniele Massaro for useful discussions about SIMD support in CUDACPP and for his help in the integration of this work in new production releases of the CUDACPP software. I warmly thank Domenico Giordano for many useful discussions about hardware benchmarking. I am grateful to the organisers and our mentors at the CSCS GPU Hackathon in Lugano in September 2022, and I thank the other participants from our team, Stephan Hageböck, Stefan Roiser and Zenny Wettersten, for our useful discussions. I kindly acknowledge the use of LUMI HPC resources under project 465001592 (CERN / Madgraph GPU porting, EHPC-DEV-2024D11-007) to derive the results I showed for AMD GPUs.